# Language Models, Graph Searching, and Supervision Adulteration: When More Supervision is Less and How to Make More More

### v1

@@ A.4 Arm Reconstruction Scratchpads Results
A.5 Graph Reconstruction Scratchpads Results
Given the path-star graph in Fig. 12(a), Figs., 12(b), 12(d), 12(c), and 12(e) illustrate the tokenization of the graph reconstruction scratchpads for all four combinations of leading-to-target or target-to-leading and either sorting by leading or target node values.
@@ A.5 Graph Reconstruction Scratchpads Results
(a) The path-star graph with $D=3$ and $M=5$ used when constructing the graph reconstruction scratchpads in Figs 12(b), 12(d), 12(c), and 12(e).
@@
Tbl. 11 shows the results for graph reconstruction scratchpads. This shows that only 4 trials succeeded in learning the task (again, ‘Test-Gen $R_{t}$’ is the statistic to consider for success).
Results for the experiments where $D=4$ and $M=5$ are plotted across training in Figs. 13(a), 13(b), 13(c), and 13(d). Figs. 13(a) and 13(b) demonstrate that sorting by leading node leads to all leading nodes being correctly predicted, regardless of whether the leading node precedes or succeeds the corresponding target node. Figs. 13(c) and 13(d) show the inverse of this where sorting on the target node leads to all target nodes being correctly predicted. Here we see a consistent pattern where the corresponding non-sorted nodes fail to be predicted above chance. Note how once the model conditions on one of these nodes, it removes it from consideration in the next prediction, hence the first one fails at $41/D=1/$ 4 chance, and the next at $31/(D-1)=1/$ 3 chance etc.
@@
A.6 Query Results
@@
Tbl. 12 shows the results of using general queries. Only 3/20 experiments succeeded in learning the task with a general target in the original setting (which uses $|V|$=100, Offline Training, and placing $Q$ after $G$) compared 16/20 corresponding experiments in the new setting (which uses $|V|=|G|$, online Training, and placing $Q$ before $G$). This indicates that such a finding may be easy to miss. However, as discussed above there is also the issue of hyperparameter-tuning a model which fails to learn a task. One can not hyperparameter tune models on the unadulterated path-star task in its original form as it doesn’t learn above chance. Thus another explanation for why this result might be hard to find, is that, without first finding working models (in our case using the causal-wise shuffling), we may not have properly set the hyperparameters needed for finding successful trials.
A.7 Tree Results
Tbl. 13 shows the results of the tree experiments. During training, we intermix sampling trees and path-star graphs with the latter being 10% of the training examples. This was done due to a length issue where trees can only make strictly shorter paths than the original path task.
@@ A.7 Tree Results
A.8 Training on Multiple Lengths and/or Degrees Results
Tbl. 14 shows the results for training using a sampled $M$ and/or sampled $D$. All values are uniformly sampled.
@@ A.8 Training on Multiple Lengths and/or Degrees Results
Appendix B Related Work
There are extensive prior works given that, a), the path-star task questions the fundamental sufficiency of next-token prediction for planning tasks, b), the presented solutions vary widely in terms of methodology, and, c), we provide theoretical insights into the task. As such, this is not an exhaustive review (and still reads like ‘A House of Leaves’ (Danielewski, 2000)) [^15] See ‘The House is Turing Complete Under Assumptions’ in Transactions of House Mathematics, 2035). To be slightly more serious, this bloated related section was done for our thesis and it would be nice if more people than just our committee ever saw it, so we included it in case others find it useful.. We also point the reader to the substantial review given in Bachmann and Nagarajan (2024).
B.1 Large Language Model (LLMs)
@@
Methods like chain-of-thought (CoT) and scratchpads elicit LLMs to generate multiple reasoning steps before generating an answer (Nye et al., 2022; Wei et al., 2022; Chu et al., 2024). This is achieved via additional prompt supervision. CoT can be done in the zero-shot setting with the generic prompt ‘Let’s think step by step’ (Kojima et al., 2022). For graphical tasks, zero-shot prompts include ‘Let’s construct a graph with the nodes and edges first’ and ‘We can use a Depth-First Search (DFS) algorithm’ Wang et al. (2023b). In the few-shot setting, prompts can provide a step-by-step decomposition of the task. For example ‘Let’s run depth-first search (DFS) step by step. Visit node 0. Neighors of node 0: [3, 6]. Visit node 6. Neighors of node 6: [3, 0]. …’ (sic., Luo et al., 2024). [^16] CoT and scratchpads use similar methods and were introduced simultaneously. ‘CoT’ is most commonly used in the training-free setting, whereas ‘scratchpads’ generally implies a training setting and supervised decomposition.
@@
B.1.1 Reasoning and Planning
While LLMs were originally designed for use on natural language tasks, it has become common to use LLMs as general predictive computation models and to apply them to reasoning tasks (Huang and Chang, 2023; Bubeck et al., 2023; Zhao et al., 2023; OpenAI, 2024), including math (Rabe et al., 2021; Zhang et al., 2022), puzzles (Shah et al., 2024; Stechly et al., 2025), code generation (Zan et al., 2023; Jiang et al., 2024b), question answering (Geva et al., 2021; Kamalloo et al., 2023; Ding et al., 2024), abstract pattern matching (Chollet et al., 2024; Chollet, 2024), graphs (see Appx. B.1.2), and planning (Zhao et al., 2023; Valmeekam et al., 2023b; Plaat et al., 2024; Stechly et al., 2025; Kang et al., 2024). Planning and reasoning are closely linked, with planning being a kind of reasoning that achieves a desired goal after a series of actions thus requiring sequential decision-making (Kang et al., 2024).
@@ B.1.1 Reasoning and Planning
We also believe that the kinds of adulteration we have described would have a small impact on the above LLMs. However, any symbolic tasks where next-tokens can be directly inferred via prior tokens, and are trained to do so, will be at risk of adulteration. This issue may become more common due to the recent interest in pretraining models on reasoning tasks. We discuss this further in Appx. B.1.2. It is unclear if in-context learning will induce the same kind of shortcuts like CHC as training, however, Khona et al. (2024) showed a simplicity bias for in-context learning, which they point out is related to shortcut learning (see Appx. B.3).
@@
Lin et al. (2025a) studies the effect of restricting training to just predicting ‘critical tokens’ instead of using full next-token prediction on reasoning tasks. They find that full next-token prediction works better for pertaining but restricted training can be more efficient for finetuning. Interestingly, the training procedure of the path-star task can be viewed as such a restriction since next-token prediction is only performed on the target-side. This is because next-token prediction on $G$ and $Q$ is invalid as both must be given information i.e. you can not predict the next token in the graph without first knowing the graph.
B.1.2 LLMs on Graphs
@@
These tasks can be specified in natural language (Tandon et al., 2019; Madaan et al., 2021; Saha et al., 2021; Sakaguchi et al., 2021; Huang et al., 2022; Valmeekam et al., 2023b; Zhang et al., 2023a; Ding et al., 2024; Huang et al., 2024a, inter alia). This introduces a subtask of mapping language to graph (Wang et al., 2023b; Fatemi et al., 2024). Madaan et al. (2022) found that LLMs that generate reasoning as code instead of natural language are better reasoners i.e. mapping to a symbolic language may offer better predictive performance.
The implicit graphical nature of reasoning tasks has motivated evaluating LLMs on explicit graphical tasks isolated from various confounding complexities that these reasoning tasks often introduce. This assumes that the minimized graphical tasks act as a surrogate to the original reasoning tasks and that this isolates aspects that make the original tasks difficult without introducing new difficulties.. To this end, many graph benchmarks and datasets have recently been introduced using synthetic data (Wang et al., 2023b; Liu and Wu, 2023; Fatemi et al., 2024; Luo et al., 2024; Chen et al., 2024b; Dai et al., 2024b, a; Fan et al., 2024) and real-world data (Guo et al., 2023; Wang et al., 2024a; Zhang et al., 2024; Wu et al., 2024a; Yuan et al., 2024; Li et al., 2024b; Tang et al., 2025). Tang et al. (2025); Fan et al. (2024) group the task by difficulty according to its complexity class (which relates to expressibility, Appx. B.5).
@@
Various things have been attributed to this poor performance. Fatemi et al. (2024) demonstrated that the way the graph is encoded in natural language for the LLM has a large impact on performance. Ge et al. (2024) found that this can be alleviated by pre-processing using some determinist ordering such as depth- or breadth-first-search. Yuan et al. (2024) found similar results with a random ordering of the graphs and showed that sorting can help. (i.e. that order matters, Appx. B.9)
Dai et al. (2024b) showed how task difficulty does not just scale with graph size but also the topology of graphs being evaluated. The path-star task is a powerful example of this, where the type of graph makes it very difficult even at small sizes, however, this isn’t an inherent property of the topology but a pathological relation between topology and training method. They also found that LLMs may apply different algorithms to various tasks and that this is sensitive to input, indicating that the LLM may be using shortcuts. Other works have also identified spurious correlations as an issue (Wang et al., 2023b, see Appx. B.3). Fatemi et al. (2024) evaluated the performance of LLMs on various graph tasks on star-shaped graphs. They found that a) the topology of graph strongly affects performance, and, b) LLMs generally do better on star-shape graphs than other types of graphs. Hallucinations have also been found to be an issue that relates to model scale and graph scale (Tang et al., 2025).
Various methods have been proposed to improve graphical reasoning: graph-specific zero-shot CoT prompts (Wang et al., 2023b, described in Appx. B.1), alternative algorithmic prompts (Dai et al., 2024b), self-prompting (Guo et al., 2023), soft-prompting (Perozzi et al., 2024), instruction-tuning (Chen et al., 2024b; Wang et al., 2024a) and instruction-tuning in conjunction with masking (Luo et al., 2024, see Appx. B.4), preference alignment, (Zhang et al., 2024; Wang et al., 2024a; Chen et al., 2024b), and re-framing the task as code for code-aware LLMs (Zhang et al., 2024; Wu et al., 2024a), which has been shown to help for other reasoning tasks (Madaan et al., 2022).
@@
Out-of-domain evaluation has also been considered. Luo et al. (2024) introduced GraphInstruct, which contains 21 graph-based tasks with 4 tasks being reserved as out-of-domain tasks that are not included in fine-tuning. Each in-domain task has 800 training examples. They used three different graph generation processes. Zhang et al. (2024) introduced NLGIFT, which included out-of-domain testing. This includes an experimental setup for fine-tuning on synthetic data and testing on real-world data. They used two different graph generation processes for the syntactic data. It has been shown that graph construction has a large impact on learnability (Saparov et al., 2025, see Appx. B.2).
We believe our work has several implications for graph benchmarks of LLMs. These works and ours have different goals and hence different research questions; these they are asking ‘how well do pretrained LLMs perform on a suit of graphical tasks?’ and then often with the secondary questions ‘why do they struggle to perform well?’ and ‘how can we improve performance post-hoc?’, whereas we are asking ‘why is learning graphical tasks hard?’. [^19] This is under the assumption that the path-star task is a minimal example of search, however, as we found, task-specific issues contribute to its difficulty. The former concerns performance while the latter is a question of learnability. From these stems the question: can the poor performance of LLMs on graphs be attributable to the same difficulties that hinder learning the path-star task? As mentioned above, we believe that adulteration will have a small impact on LLMs. However, the issues we present will become more applicable as people move to pretraining LLMs for reasoning tasks. These issues may also affect finetuning. Thus our work motivates the careful design of graph tasks when training or finetuning models. We leave it to future work to see if our methods can be used to improve the performance of LLMs.
@@
B.2 Learnability of Graphs
Unlike the above works evaluating LLMs, we are concerned with the learnability of graph algorithms on decoder-only (transformer) language models. Saparov et al. (2025) is the most closely related work (outside of Bachmann and Nagarajan (2024)). They consider finding the shortest path given a graph. As with our experimental setup, they provide a query with a start and end node, and the graph is encoded as a list of shuffled edges. The graph is also randomly generated and is semanticless.
Their first finding is that graph topology highly affects performance (especially in out-of-domain evaluation across topologies). This was also observed by Dai et al. (2024b). They find that a ‘balanced’ graph topology works the best. These are graphs sampled from a generative process which creates a graph with a uniform distribution over the number of ‘lookaheads’ (path length) required to solve the task. We conjecture that these work well because they better support task decomposition.
As our work was nearly completed before we became aware of their work, we do not do direct comparisons. There are several differences: 1) Most of their experiments use encoder-only models. They did not evaluate path-star graphs using decoder-only models (only using encoder models, like Frydenlund (2024)). 2) They employed a slight architecture modifications that concatenates the token and position embeddings. 3) They used rotary positional embeddings in their decoder-only experiments (again, only on balanced graphs). 4) They used an approximate second-order optimizer, Sophia (where we used Adam). 5) Their best models were also trained for 883M samples (where we used 100M). We suspect that all of these may contribute to differences in performance.

### v3

@@ B.4 Arm Reconstruction Scratchpads Results
B.5 Graph Reconstruction Scratchpads Results
Given the path-star graph in Fig. 21(a), Figs., 21(b), 21(d), 21(c), and 21(e) illustrate the tokenization of the graph reconstruction scratchpads for all four combinations of leading-to-target or target-to-leading and either sorting by leading or target node values.
@@ B.5 Graph Reconstruction Scratchpads Results
(a) The path-star graph with $D=3$ and $M=5$ used when constructing the graph reconstruction scratchpads in Figs 21(b), 21(d), 21(c), and 21(e).
@@
Tbl. 6 shows the results for graph reconstruction scratchpads. This shows that only 4 trials succeeded in learning the task (again, ‘Test-Gen $R_{t}$’ is the statistic to consider for success).
Results for the experiments where $D=4$ and $M=5$ are plotted across training in Figs. 22(a), 22(b), 22(c), and 22(d). Figs. 22(a) and 22(b) demonstrate that sorting by leading node leads to all leading nodes being correctly predicted, regardless of whether the leading node precedes or succeeds the corresponding target node. Figs. 22(c) and 22(d) show the inverse of this where sorting on the target node leads to all target nodes being correctly predicted. Here we see a consistent pattern where the corresponding non-sorted nodes fail to be predicted above chance. Note how once the model conditions on one of these nodes, it removes it from consideration in the next prediction, hence the first one fails at $41/D=1/$ 4 chance, and the next at $31/(D-1)=1/$ 3 chance etc.
@@
Note that had we given the full arms instead of just pairs of nodes as the scratchpad, the task should become learnale, though for incorrect reasons, where instead of learning planning or arm reconstruction, the model should be able to use the sorting shortcuts in conjunction with the CHC to get 100% accuracy. That is, the version we present here is designed to avoid adulteration, even if that would actually lead to ‘succeeding’ on the task.
B.6 Query Results
@@
Tbl. 7 shows the results of using general queries. Only 3/20 trials succeeded in learning the task with a general target in the original setting (which uses $|V|$=100, Offline Training, and placing $Q$ after $G$) compared 16/20 successful corresponding trials in the new setting (which uses $|V|=|G|$, online Training, and placing $Q$ before $G$). This indicates that such a finding may be easy to miss. However, as discussed above there is also the issue of hyperparameter-tuning a model which fails to learn a task. One can not hyperparameter tune models on the unadulterated path-star task in its original form as it doesn’t learn above chance. Thus another explanation for why this result might be hard to find, is that, without first finding working models (in our case using the causal-wise shuffling), we may not have properly set the hyperparameters needed for finding successful trials.
B.7 Tree Results
Tbl. 8 shows the results of the tree experiments. During training, we intermix sampling trees and path-star graphs with the latter being 10% of the training examples. This was done due to a length issue where trees can only make strictly shorter paths than the original path task.
@@ B.7 Tree Results
B.8 Training on Multiple Lengths and/or Degrees Results
Tbl. 9 shows the results for training using a sampled $M$ and/or sampled $D$. All values are uniformly sampled.
@@ B.8 Training on Multiple Lengths and/or Degrees Results
Appendix C Related Work
There are extensive prior works given that, a), the path-star task questions the fundamental sufficiency of next-token prediction for planning tasks, b), the presented solutions vary widely in terms of methodology, and, c), we provide theoretical insights into the task. As such, this is not an exhaustive review (and still reads like ‘A House of Leaves’ (Danielewski, 2000)) [^24] See ‘The House Is Turing Complete Under Assumptions’ in Transactions of House Mathematics, 2035). To be slightly more serious, this bloated related works was done for our thesis and it would be nice if more people than just our committee ever saw it, so we included it in case others find it useful.. We also point the reader to the substantial review given in Bachmann and Nagarajan (2024).
C.1 Large Language Model (LLMs)
@@
Methods like chain-of-thought (CoT) and scratchpads elicit LLMs to generate multiple reasoning steps before generating an answer (Nye et al., 2022; Wei et al., 2022; Chu et al., 2024). This is achieved via additional prompt supervision. CoT can be done in the zero-shot setting with the generic prompt ‘Let’s think step by step’ (Kojima et al., 2022). For graphical tasks, zero-shot prompts include ‘Let’s construct a graph with the nodes and edges first’ and ‘We can use a Depth-First Search (DFS) algorithm’ (Wang et al., 2023b). In the few-shot setting, prompts can provide a step-by-step decomposition of the task. For example ‘Let’s run depth-first search (DFS) step by step. Visit node 0. Neighors of node 0: [3, 6]. Visit node 6. Neighors of node 6: [3, 0]. …’ (sic., Luo et al., 2024). [^25] CoT and scratchpads use similar methods and were introduced simultaneously. ‘CoT’ is most commonly used in the training-free setting, whereas ‘scratchpads’ generally implies a training setting and supervised decomposition.
@@
C.1.1 Reasoning and Planning
While LLMs were originally designed for use on natural language tasks, it has become common to use LLMs as general predictive computation models and to apply them to reasoning tasks (Huang and Chang, 2023; Bubeck et al., 2023; Zhao et al., 2023; OpenAI, 2024), including math (Rabe et al., 2021; Zhang et al., 2022), puzzles (Shah et al., 2024; Stechly et al., 2025), code generation (Zan et al., 2023; Jiang et al., 2024b), question answering (Geva et al., 2021; Kamalloo et al., 2023; Ding et al., 2024), abstract pattern matching (Chollet et al., 2024; Chollet, 2024), graphs (see Appx. C.1.2), and planning (Zhao et al., 2023; Valmeekam et al., 2023b; Plaat et al., 2024; Stechly et al., 2025; Kang et al., 2024). Planning and reasoning are closely linked, with planning being a kind of reasoning that achieves a desired goal after a series of actions thus requiring sequential decision-making (Kang et al., 2024).
@@ C.1.1 Reasoning and Planning
We also believe that the kinds of adulteration we have described would have a small impact on the above LLMs. However, any symbolic tasks where next-tokens can be directly inferred via prior tokens, and are trained to do so, will be at risk of adulteration. This issue may become more common due to the recent interest in pretraining models on reasoning tasks. We discuss this further in Appx. C.1.2. It is unclear if in-context learning will induce the same kind of shortcuts like CHC as training, however, Khona et al. (2024) showed a simplicity bias for in-context learning, which they point out is related to shortcut learning (see Appx. C.3).
@@
Lin et al. (2025a) studies the effect of restricting training to just predicting ‘critical tokens’ instead of using full next-token prediction on reasoning tasks. They find that full next-token prediction works better for pertaining but restricted training can be more efficient for finetuning. Interestingly, the training procedure of the path-star task can be viewed as such a restriction since next-token prediction is only performed on the target-side. This is because next-token prediction on $G$ and $Q$ is invalid as both must be given information, i.e., you can not predict the next token in the graph without first knowing the graph.
C.1.2 LLMs on Graphs
@@
These tasks can be specified in natural language (Tandon et al., 2019; Madaan et al., 2021; Saha et al., 2021; Sakaguchi et al., 2021; Huang et al., 2022; Valmeekam et al., 2023b; Zhang et al., 2023a; Ding et al., 2024; Huang et al., 2024a, inter alia). This introduces a subtask of mapping language to graph (Wang et al., 2023b; Fatemi et al., 2024). Madaan et al. (2022) found that LLMs that generate reasoning as code instead of natural language are better reasoners, i.e., mapping to a symbolic language may offer better predictive performance.
The implicit graphical nature of reasoning tasks has motivated evaluating LLMs on explicit graphical tasks isolated from various confounding complexities that these reasoning tasks often introduce. This assumes that the minimized graphical tasks act as a surrogate to the original reasoning tasks and that this isolates aspects that make the original tasks difficult without introducing new difficulties.. To this end, many graph benchmarks and datasets have recently been introduced using synthetic data (Wang et al., 2023b; Liu and Wu, 2023; Fatemi et al., 2024; Luo et al., 2024; Chen et al., 2024b; Dai et al., 2024b, a; Fan et al., 2024) and real-world data (Guo et al., 2023; Wang et al., 2024a; Zhang et al., 2024; Wu et al., 2024a; Yuan et al., 2024; Li et al., 2024b; Tang et al., 2025). Tang et al. (2025); Fan et al. (2024) group the task by difficulty according to its complexity class (which relates to expressibility, Appx. C.5).
@@
Various things have been attributed to this poor performance. Fatemi et al. (2024) demonstrated that the way the graph is encoded in natural language for the LLM has a large impact on performance. Ge et al. (2024) found that this can be alleviated by pre-processing using some determinist ordering such as depth- or breadth-first-search. Yuan et al. (2024) found similar results with a random ordering of the graphs and showed that sorting can help. (i.e., that order matters, Appx. C.9)
Dai et al. (2024b) showed how task difficulty does not just scale with graph size but also the topology of graphs being evaluated. The path-star task is a powerful example of this, where the type of graph makes it very difficult even at small sizes, however, this isn’t an inherent property of the topology but a pathological relation between topology and training method. They also found that LLMs may apply different algorithms to various tasks and that this is sensitive to input, indicating that the LLM may be using shortcuts. Other works have also identified spurious correlations as an issue (Wang et al., 2023b, see Appx. C.3). Fatemi et al. (2024) evaluated the performance of LLMs on various graph tasks on star-shaped graphs. They found that a) the topology of graph strongly affects performance, and, b) LLMs generally do better on star-shape graphs than other types of graphs. Hallucinations have also been found to be an issue that relates to model scale and graph scale (Tang et al., 2025).
Various methods have been proposed to improve graphical reasoning: graph-specific zero-shot CoT prompts (Wang et al., 2023b, described in Appx. C.1), alternative algorithmic prompts (Dai et al., 2024b), self-prompting (Guo et al., 2023), soft-prompting (Perozzi et al., 2024), instruction-tuning (Chen et al., 2024b; Wang et al., 2024a) and instruction-tuning in conjunction with masking (Luo et al., 2024, see Appx. C.4), preference alignment, (Zhang et al., 2024; Wang et al., 2024a; Chen et al., 2024b), and re-framing the task as code for code-aware LLMs (Zhang et al., 2024; Wu et al., 2024a), which has been shown to help for other reasoning tasks (Madaan et al., 2022).
@@
Out-of-domain evaluation has also been considered. Luo et al. (2024) introduced GraphInstruct, which contains 21 graph-based tasks with 4 tasks being reserved as out-of-domain tasks that are not included in fine-tuning. Each in-domain task has 800 training examples. They used three different graph generation processes. Zhang et al. (2024) introduced NLGIFT, which included out-of-domain testing. This includes an experimental setup for fine-tuning on synthetic data and testing on real-world data. They used two different graph generation processes for the syntactic data. It has been shown that graph construction has a large impact on learnability (Saparov et al., 2025, see Appx. C.2).
We believe our work has several implications for graph benchmarks of LLMs. These works and ours have different goals and hence different research questions; these they are asking ‘how well do pretrained LLMs perform on a suit of graphical tasks?’ and then often with the secondary questions ‘why do they struggle to perform well?’ and ‘how can we improve performance post-hoc?’, whereas we are asking ‘why is learning graphical tasks hard?’. [^28] This is under the assumption that the path-star task is a minimal example of search on graphs, however, as we found, task-specific issues contribute to its difficulty. The former concerns performance while the latter is a question of learnability. From these stems the question: can the poor performance of LLMs on graphs be attributable to the same difficulties that hinder learning the path-star task? As mentioned above, we believe that adulteration will have a small impact on LLMs. However, the issues we present will become more applicable as people move to pretraining LLMs for reasoning tasks. These issues may also affect finetuning. Thus our work motivates the careful design of graph tasks when training or finetuning models. We leave it to future work to see if our methods can be used to improve the performance of LLMs.
@@
C.2 Learnability of Graphs
Unlike the above works evaluating LLMs, we are concerned with the learnability of graph algorithms on decoder-only transformer language models. Saparov et al. (2025) is the most closely related work (outside of Bachmann and Nagarajan (2024)). They consider finding the shortest path given a graph. As with our experimental setup, they provide a query with a start and end node, and the graph is encoded as a list of shuffled edges. The graph is also randomly generated and semanticless.
Their first finding is that graph topology highly affects performance (especially in out-of-domain evaluation across topologies). This was also observed by Dai et al. (2024b). They find that a ‘balanced’ graph topology works the best. These are graphs sampled from a generative process which creates a graph with a uniform distribution over the number of ‘lookaheads’ (path length) required to solve the task. Note that, while we discussed these as ‘more general graphs’ in the main text, they are very closely related to path-star graphs.
As our work was nearly completed before we became aware of their work, we do not do direct comparisons. There are several differences: 1) Most of their experiments use encoder-only models. They did not evaluate path-star graphs using decoder-only models (only using encoder-only models, as in Frydenlund (2024)). 2) They employed a slight architecture modifications that concatenates the token and position embeddings. 3) They used rotary positional embeddings in their decoder-only experiments (again, only on balanced graphs). 4) They used an approximate second-order optimizer, Sophia (where we used Adam). 5) Their best models were also trained for 883M samples (where we used 100M). We suspect that all of these may contribute to differences in performance.